OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (2, 4) and a directrix of y = 8.

OpenStudy (anonymous):

f(x) = −1/8 (x − 2)2 + 6 f(x) = 1/8 (x − 2)2 + 6 f(x) = −1/8 (x + 2)2 + 8 f(x) = 1/8 (x + 2)2 + 8

OpenStudy (anonymous):

@ganeshie8 @Haseeb96

OpenStudy (anonymous):

look The equation of a parabola with focus in (h,k+p) and directrix y=k-p, is, (x−h)2=4p(y−k) can you do the rest ?

OpenStudy (anonymous):

whats h and k and p?

OpenStudy (anonymous):

@pak

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@Brainybeauty

OpenStudy (anonymous):

I'm sorry I have not yet studied parabolas... :/

OpenStudy (anonymous):

Oh okay, thanks for trying though @Brainybeauty

OpenStudy (paki):

do you know the basics of parabola...?

OpenStudy (anonymous):

:) Ur welcome. Do tag me in your other questions. I will try to solve them if I can! :)

OpenStudy (anonymous):

No I don't @paki

OpenStudy (anonymous):

f(x) = 1/8 (x + 2)2 + 8 ?

OpenStudy (paki):

how... can you explain...?

OpenStudy (solomonzelman):

This is not going anywhere, if you try to derive the equation of the parabola without knowing much about the concept "parabola".

OpenStudy (anonymous):

Parabolas are because of quadratic equations. The values which solve quadratic equations make a parabola.

OpenStudy (solomonzelman):

Yes, if the parabola opens up or down.

OpenStudy (anonymous):

Yes. If x is negative, the parabola is a straight U and if it is positive it is an inverted U.

OpenStudy (anonymous):

I gave up on the question. The joke was what type of shoes do ninjas wear? Sneakers @Brainybeauty @paki

OpenStudy (anonymous):

Haha! That's funny XD

OpenStudy (paki):

lol.... HEELS :)

OpenStudy (anonymous):

But please try you will get it. The value of x should be such that f(x) is equal to 8.

OpenStudy (anonymous):

bacuase directrix of y=8 and y is usually the function of x so...

OpenStudy (solomonzelman):

Brainybeauty, it is the other way, if ax²+bx+c=0 is with a positive a, then it opens up, and if a is a negative number than it opens down (and if a=0 then it is a line).

OpenStudy (anonymous):

I don't know much about this topic, it hasn't been discussed in my school yet, but I'm just recalling what I heard some people say...

OpenStudy (solomonzelman):

Yes, but it is good to remember, a>0 opens up a=0 line a<0 opens down

OpenStudy (solomonzelman):

Horizontal parabola is not a function.

OpenStudy (anonymous):

Yesss! I get it now. Thx. :)

OpenStudy (solomonzelman):

Anytime:) I myself a trig student. I know some things from calc 1 though....

OpenStudy (anonymous):

Wow! I'm studying trigo too (basic and its difficulty is increasing). But I don't know anything about calculus. I just know bits about differentiation and limits. I am going to study it next year. @SolomonZelman